OpenStudy (anonymous):

Will Fan and Medal! This table shows the weights of four boxes that need to be mailed. Which choice lists the boxes ordered from heaviest to lightest? Box Weight box 1 2.7 oz box 2 1.3 lb box 3 2.6 lb box 4 1.9 oz A.box 1, box 3, box 4, box 2 B.box 3, box 2, box 1, box 4 C.box 1, box 2, box 3, box 4 D.box 4, box 1, box 2, box 3

OpenStudy (anonymous):

d ounces weigh less than pounds

OpenStudy (anonymous):

Okay! Ill see if its right and ill get back to you :)

OpenStudy (doc.brown):

\[1lb = 16oz\]

OpenStudy (anonymous):

It was incorrect :c The Answer is B

OpenStudy (doc.brown):

\[box_2=1.3lb\times\frac{16oz}{lb}\]\[box_2=1.3\cancel{lb}\times\frac{16oz}{\cancel{lb}}\]\[box_2=1.3\times16oz\]\[box_2=20.8oz\]

OpenStudy (anonymous):

sorry i did lightest to heaviest